OpenStudy (anonymous):

13x^2/x+4-16+12x^2/x+4 subtract the rational expressions and simplify if possible

jhonyy9 (jhonyy9):

13x2 16+12x2 13x2 - 16 - 12x2 x2 -16 (x-4)(x+4) ------ - --------- = ----------------- = -------- = ----------- = x+4 x+4 x+4 x+4 x+4 = x-4